# Bayesian reconstruction of primordial perturbations from induced gravitational waves

### v2

@@ 5 Conclusions
Our method can be extended and applied to more general situations. It would be interesting to include the effects of primordial curvature non-Gaussianities which can influence the SIGW profile (see e.g. [66, 67, 68, 69]). Besides curvature fluctuations, also isocurvature modes might play a role in SIGW generation (see e.g.  [70, 71, 72, 73, 74, 75]). Although we focussed on scalar sources, primordial tensor degrees of freedom can induce GW as well [76]. Besides considering more general primordial sources, it would also be interesting to model more accurately our priors regarding the infrared parts of the curvature and GW spectra; to include accurate constraints related to the corresponding primordial black hole formation; and to analyse how imperfect knowledge of instrumental noise might affect our results. Finally, suitable extensions of our method can allow not only to infer the $\mathcal{P}_{\zeta}$ profile, but also to reconstruct the inflationary scalar potential leading to SIGW profiles. We hope to return soon to these topics in separate publications.
@@ Acknowledgements
We thank Sukannya Bhattacharya, Jonas El Gammal, Gabriele Franciolini, Juan Garcia-Bellido, Jan Hamann, Marco Peloso, and David Wands for useful discussions. We are partially funded by the STFC grants ST/T000813/1 and ST/X000648/1. A.G. is supported by the UKRI AIMLAC CDT, funded by grant EP/S023992/1. We also acknowledge the support of the Supercomputing Wales project, which is part-funded by the European Regional Development Fund (ERDF) via Welsh Government. For the purpose of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Research Data Access Statement: No new data were generated for this manuscript.
@@ Appendix A Scalar power spectrum reconstruction with recent PTA results
We test our method against real data from Pulsar Timing Arrays.131313 See e.g. [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] for other studies. We note that the scalar induced GW interpretation of the PTA results is likely ruled out due to overproduction of PBH, at least in the case of Gaussian statistics of $\zeta$ and assuming radiation domination. The inclusion of non-Gaussianity or a non-standard equation of state may alleviate this, as studied in refs [89, 86, 90, 88, 87, 91, 92, 93]. We make use of the public package Ceffyl [94] to interface our code with the NANOGrav 15 year data [95]. The results of the reconstruction are plotted in fig. 12. We do not find any evidence for significant deviations from a power-law $\mathcal{P}_{\zeta}$ or $\Omega_{\mathrm{GW}}$ with little to choose amongst the different models, with the $\log\mathcal{Z}=-63.71,-63.66-63.60,-63.59,-64.57$ for $N=2,3,4,5,6$ respectively.
Figure 12: Spline based reconstruction of the curvature power spectrum from PTA (NANOGrav 15 year). Left: The power spectrum posterior Right: The corresponding $\Omega_{\rm GW}$ spectrum along with violin plots of the free spectrum HD reconstruction from NG15.

### v3

@@ 5 Conclusions
Our method can be extended and applied to more general situations. It would be interesting to include the effects of primordial curvature non-Gaussianities which can influence the SIGW profile (see e.g. [Unal:2018yaa, Cai:2018dig, Adshead:2021hnm, Perna:2024ehx]). Besides curvature fluctuations, also isocurvature modes might play a role in SIGW generation (see e.g.  [Domenech:2023jve, Passaglia:2021jla, Domenech:2021and, Marriott-Best:2025sez, Ozsoy:2023gnl, Chen:2024twp]). Although we focussed on scalar sources, primordial tensor degrees of freedom can induce GW as well [Bari:2023rcw]. Besides considering more general primordial sources, it would also be interesting to model more accurately our priors regarding the infrared parts of the curvature and GW spectra; to include accurate constraints related to the corresponding primordial black hole formation; and to analyse how imperfect knowledge of instrumental noise might affect our results. Finally, suitable extensions of our method can allow not only to infer the $\mathcal{P}_{\zeta}$ profile, but also to reconstruct the inflationary scalar potential leading to SIGW profiles. We hope to return soon to these topics in separate publications.
@@ Acknowledgements
We thank Sukannya Bhattacharya, Jonas El Gammal, Gabriele Franciolini, Juan Garcia-Bellido, Jan Hamann, Marco Peloso, and David Wands for useful discussions. We are partially funded by the STFC grants ST/T000813/1 and ST/X000648/1. A.G. is supported by the UKRI AIMLAC CDT, funded by grant EP/S023992/1. We also acknowledge the support of the Supercomputing Wales project, which is part-funded by the European Regional Development Fund (ERDF) via Welsh Government. For the purpose of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Research Data Access Statement: The code used to generate the mock data and run the analysis is available at [Ameek94SIGWInverse].
@@ Appendix A Scalar power spectrum reconstruction with recent PTA results
We test our method against real data from Pulsar Timing Arrays. [^14] See e.g. [NANOGrav:2023hvm, EPTA:2023xxk, Figueroa:2023zhu, Ellis:2023oxs, Cai:2023dls, Yi:2023mbm, Yi:2023tdk, Firouzjahi:2023lzg, Liu:2023pau, Wang:2023ost, Domenech:2024rks, Harigaya:2023pmw] for other studies. We note that the scalar induced GW interpretation of the PTA results is likely ruled out due to overproduction of PBH, at least in the case of Gaussian statistics of $\zeta$ and assuming radiation domination. The inclusion of non-Gaussianity or a non-standard equation of state may alleviate this, as studied in refs [Balaji:2023ehk, Wang:2023ost, Franciolini:2023pbf, Harigaya:2023pmw, Domenech:2024rks, Liu:2023ymk, Liu:2023hpw, Inui:2024fgk]. We make use of the public package Ceffyl [lamb2023rapid] to interface our code with the NANOGrav 15 year data [NANOGrav:2023gor, the_nanograv_collaboration_2025_16051178]. The results of the reconstruction are plotted in fig. 12. We do not find any evidence for significant deviations from a power-law $\mathcal{P}_{\zeta}$ or $\Omega_{\mathrm{GW}}$ with little to choose amongst the different models, with the $\log\mathcal{Z}=-63.71,-63.66-63.60,-63.59,-64.57$ for $N=2,3,4,5,6$ respectively.
Figure 12: Spline based reconstruction of the curvature power spectrum from PTA (NANOGrav 15 year). Left: The power spectrum posterior Right: The corresponding $\Omega_{\rm GW}$ spectrum along with violin plots of the free spectrum Hellings Downs reconstruction from NG15.
@@ Appendix B General equation of state and broad spectra
Appendix C Noisy realisations of GW data
To test the robustness of our method, we also apply it to noisy realisations of the mean spectra presented in section 4.1 for which we generate random realisations drawn from a Gaussian distribution defined by eq. 4.2. Our method can again recover the a curvature power spectrum close to the actual one, as long as the noise level close to the peak of the GW spectrum is low i.e. the shape of the peak of the noisy realisation is not too different from the mean spectrum. As discussed earlier, this is a consequence of the fact that the shape of the induced GW spectrum depends most strongly on the shape of the peak of the curvature power spectrum. The latter itself gets tightly constrained by the low noise levels around the peak of the GW spectrum.
The Bayesian evidence shows a clear trend for the BPL model, irrespective of the individual realisation, with models with more than 6 nodes being favoured less, as shown in Figure 15. For brevity, we only show the reconstruction for one particular realisation. We also plot the fidelity of the reconstructed $P_{\zeta}$ and $\Omega_{\rm GW}$ for this realisation, showing $\Delta P_{\zeta}\equiv|\hat{P}_{\zeta}-P_{\zeta}^{\rm true}|/P_{\zeta}^{\rm true}$ where $\hat{P}_{\zeta}$ represents the median of the reconstruction and similarly for $\Omega_{\rm GW}$. We see that the relative error remains small, reaching $10^{-1}$ or lower near the peak and only gets higher far away from it, especially in the IR for $P_{\zeta}$.
Figure 15: Top: Reconstruction of the BPL spectrum for a particular noisy realisation of the GW data using linear interpolation. Bottom: $\log\mathcal{Z}$ values across five different realisations.
For the peaked and oscillatory spectra we find that we are still able to accurately recover the underlying curvature power spectrum. However, increasing the number of nodes does not immediately lead to a decrease of the evidence independently of the random realisation, owing to the increased complexity of the models, similar to what we observed with the mean-only realisations in section 4.1.